# citations.bib

### v2

@@ I Introduction
High-performance computing (HPC) represents the pinnacle of computational power, harnessing clusters of computing resources to transcend the limitations of individual machines. The fundamental advantage of HPC lies in its implementation of parallel processing techniques that maximize the performance of processor clusters, enabling complex data applications and mathematical calculations that would otherwise be infeasible [Nichols_2024]. HPC has been instrumental in driving innovation across diverse domains including climate modeling, astrophysics simulations, pharmaceutical research, energy optimization, financial risk analysis, and training state-of-the-art machine learning models, particularly Large Language Models [chen2024landscapechallengeshpcresearch, ding2023hpc, hoffmann2022training, jin2023evaluation, narayanan2021efficient, yin2023ion]
@@
Large Language Models (LLMs) have revolutionized technology across multiple domains, democratizing access to high-level code through natural language interfaces. These models have evolved from conversational assistants to autonomous entities capable of executing tasks traditionally performed manually, such as code generation [jiang2024survey, wang2024llm, liu2024evaluating, phuong2024TDD, li2024using]. The ability of LLMs to generate, test, and evaluate code has significantly enhanced software engineering capabilities, reducing development cycles and increasing productivity.
@@
Despite their general utility, LLMs face significant challenges when applied to domain-specific knowledge requirements. According to the ParEval benchmark, both open-source and commercial LLMs demonstrate substantially lower accuracy in generating correct and efficient parallel code compared to serial code [Nichols_2024]. This limitation stems from LLMs’ training methodology, which typically involves diverse internet-sourced data rather than specialized domain knowledge. The relative scarcity of HPC code examples available online further exacerbates this knowledge gap [Nichols_2024, chen2024landscapechallengeshpcresearch]. LLMs have also been successfully deployed in parallel program verification [aaron_iwomp24, sollenberger2024llm4vvexploringllmasajudgevalidation]. However, the unique demands of high-performance computing—requiring both rigorously tested and optimized code for parallelism, memory efficiency, and execution speed—raise a critical research question: How can we effectively leverage LLMs to comprehensively address the HPC-specific programming challenges?
@@
A promising approach to addressing LLM limitations is the implementation of Artificial Intelligence (AI) Agents—autonomous LLM-based systems capable of performing complex tasks through customized workflows and specialized tools [xi2023survey, acharya2024agentic, li2024survey, gupta2024survey, sudalaimuthu2024enhancing]. AI agents extend beyond natural language processing to encompass decision-making, problem-solving, and execution capabilities. By utilizing external tools, these agents can access up-to-date information and effectively complete domain-specific tasks. Multi-agent systems further enhance this capability by distributing complex workloads across specialized agents, enabling the handling of tasks that exceed the capabilities of any single agent.
@@ I-A Research Contributions
  • Specialized Agent Architecture: MARCO introduces a clear separation between code generation, performance evaluation and testing agents, allowing each to develop distinct expertise. The code generation agent specializes in implementing advanced optimization techniques including cache locality optimization, parallelization through OpenMP or CUDA, and vectorization strategies. Meanwhile, the evaluation agent focuses on rigorously assessing execution time, memory usage, and algorithmic complexity, while the testing agent continuously evaluates output correctness.
@@
To assess the effectiveness of our approach, we first assembled a thorough evaluation dataset, which included the LeetCode 75 problem set along with an additional 10 challenging problems. Next, we implemented MARCO and compared its performance to leading language models such as GPT-4o, Claude 3.5 Sonnet, DeepSeek Coder-V2, and Llama 3.1. Additionally, we conducted detailed ablation studies to measure the contribution of each component within the MARCO architecture.
@@
The remainder of this paper is organized as follows: Section II discusses background and related work, Section III details our methodology, Section IV presents experimental results, Section VI provides an in-depth analysis of the MARCO system, Section VII discusses our findings, and Section VIII concludes with future research directions.
@@ II-A LLMs for Code Generation
LLMs are able to perform a wide range of tasks, such as code generation and understanding, through parameters and their transformer architecture [wang2024llm, liu2024evaluating, phuong2024TDD, li2024survey]. LLMs have billions of parameters that enable the LLM to pick up on patterns that are passed into it during the training. The relationship between parameter size and model effectiveness is linear; with an increase in the number of parameters, the chance that the LLM can solve higher-complexity or data-intensive problems increases [kaplan2020scaling, hoffmann2022training, earle2023level]. Alongside the high number of parameters is the transformer architecture, otherwise known as neural networks. These networks are intended to mimic the structure of a human brain’s neurons to predict intended outcomes.
@@
Chatbots like ChatGPT were the first to provide general-purpose assistance across numerous domains. LLMs gain knowledge of these domains through training with large amounts of data being fed into the model. OpenAI, for example, trained ChatGPT using techniques like Reinforcement Learning from Human Feedback [ouyang2022training, christiano2017deep, liu2024best], which utilizes a reward system to associate good responses with confirmation from a human trainer. With the adaptability of LLMs, specialization has occurred with the creation of new models to focus on a specific domain for designated tooling. One popular domain of specialization is coding, where the model is mostly pre-trained on code samples to improve its ability to handle coding situations. Some code LLMs like Qwen2.5-Coder utilize other techniques like Synthetic Data to deal with issues regarding the scarcity of high-level code [hui2024qwen, wang2024llm].
@@
While LLMs have demonstrated impressive performance in code generation tasks, their effectiveness is primarily confined to small-scale problems, such as specific functions and single-file programs [li2024performance]. As the complexity and scale of coding problems increase, these models experience a significant decline in performance, largely due to the limitations in their ability to retain and process larger contexts [wang2024large]. This degradation is directly related to the LLM’s context window—the maximum amount of information the model can consider when generating responses, measured in tokens. LLMs process natural language by breaking it down into tokens, which can range from individual characters to complete words, allowing them to interpret and generate human-like text [balfroid2024towards]. Research by IBM has demonstrated that context window size is a critical factor in determining an LLM’s effectiveness for complex tasks like code generation, as larger windows enable models to ingest more documentation and maintain awareness of broader code structures [antropic2023introducing]. The evolution of this capability is evident in the progression from older models like GPT-3.5, with a context window of 16,385 tokens, to more advanced models such as GPT-4.1, which boasts a significantly expanded capacity of 1,047,576 tokens [li2024performance].
@@ II-B Challenges in LLM-Generated HPC Code
LLM-generated code often exhibits significant deficiencies in performance optimization, including inefficient cache utilization and suboptimal parallelization strategies [nichols2024large]. The ParEval benchmark has demonstrated a substantial gap between LLMs’ ability to generate serial versus parallel code, with all evaluated models showing markedly lower accuracy when tasked with parallel code generation [nichols2024large, davis2024pareval]. While open-source repositories provide vast amounts of training data for coding tasks, this breadth comes at the expense of specialized domain knowledge required for high-performance computing applications [ding2023hpc]. The primary challenge stems from how LLMs process information through Natural Language Processing (NLP) techniques, which excel with textual inputs but struggle with the computationally intensive scientific data prevalent in HPC contexts [chen2023lm4hpc].
Modern HPC infrastructure leverages parallel execution on multi-core systems through threading and synchronization mechanisms to optimize code performance; however, due to the relatively recent emergence of these architectures, LLMs have limited contextual understanding of them [kadosh2023scope]. This knowledge gap results in parallel code generation that performs substantially worse than sequential implementations [nichols2024large]. Furthermore, the advanced algorithms and optimization techniques typical in HPC environments create additional complexity that makes debugging errors particularly challenging for current LLM implementations [ding2023hpc].
These contextually intense factors to keep in context make general-purpose LLMs struggle to apply all of them in success. Recent research has started in this problem space with the introduction of HPC-Coder, which utilizes fine-tuning a pre-existing model on HPC data to improve its ability to write parallel code [nichols2023hpc]. This fine-tuned model outperforms other general-purpose LLMs on HPC-specific tasks and OpenMP pragma labeling, displaying the potential for HPC optimization to LLMs.
LLM-generated code also struggles to get complex questions correct on the first try. Some LLMs may get a question correct on the third or fifth attempt [huang2023agentcoder]. LLMs are limited by this approach of trying to get something correct on the first try. An agent approach may be needed for more difficult problems and complex code. The difference with an agent is that it can have multiple attempts to get a question correct, as well as multiple attempts to improve the code that it is trying to create [dong2023selfcollab]. This is similar to a human trying a question, realizing what they did wrong, and improving their answer [yuqian2024colearning]. For limitations of LLMs in regard of HPC code generation, a simple LLM approach may not suffice [singh2023selfplanning].
@@ III-A Research Contributions
Current approaches to code generation using Large Language Models (LLMs) exhibit significant limitations when applied to high-performance computing (HPC) workloads. Most existing models generate code in a single-pass manner, focusing primarily on syntactic correctness rather than optimizing for parallelism, memory hierarchy, or the architectural characteristics of the target machine. As a result, the generated code often underperforms in terms of memory efficiency, execution speed, and scalability, especially when compared to hand-optimized or domain-specific implementations. This performance gap leaves substantial room for manual tuning and expert intervention, underscoring the inadequacy of current LLM-driven code generation workflows for HPC applications. Recent studies, such as HPC-GPT [ding2023hpc] and the ParEval benchmark [nichols2024large], have empirically demonstrated that LLMs struggle to deliver performant parallel code, revealing a clear need for more sophisticated, architecture-aware generation techniques. Furthermore, work by Nichols et al. [nichols2023hpc] shows that even fine-tuned models on HPC-specific data require additional optimization layers to approach the performance of expert-written code.
@@
To address these challenges, we introduce MARCO (Multi-Agent Reactive Code Optimizer), a multi-agent system that fundamentally reimagines the code optimization process by combining automated code generation and systematic performance testing, leveraging API integrations for dynamic adaptation. This modular separation between generation and evaluation enables improvements beyond conventional monolithic approaches. MARCO’s architecture is centered on four key innovations:
@@
  • Specialized Agent Separation: MARCO implements a clear division of responsibilities between the code optimizer and the testing agent. The optimizer agent focuses on applying advanced strategies such as cache locality improvements, parallelization through OpenMP or CUDA, and vectorization techniques, drawing upon insights from recent high-performance computing literature [ding2023hpc, nichols2023hpc]. Meanwhile, the testing agent rigorously evaluates execution time, memory consumption, algorithmic complexity, output correctness, and incremental improvements over the original unoptimized code.
@@
  • Adaptive Feedback Loop: Unlike traditional single-pass LLM systems that search for immediate solutions, MARCO employs an iterative feedback mechanism. Performance metrics gathered by the testing agent are fed back to the optimizer, creating a refinement loop that progressively drives the system toward more efficient implementations, aligning with multi-step agentic workflows recently explored in AI systems research [huang2023agentcoder].
@@
  • Cost-Efficient Design: By decoupling optimization and evaluation tasks, MARCO reduces unnecessary API calls and minimizes token consumption at each iteration. This architectural design ensures that overall system costs remain controlled, offering a scalable, resource-aware alternative to computationally expensive fine-tuning approaches.
@@
  • Broad Applicability: While many existing frameworks rely on specialized fine-tuning for narrow HPC tasks, MARCO dynamically adapts to live performance data and integrates real-time insights from web-sourced research. This generalizable design reduces the time and resources required for model retraining, maintaining efficiency across a diverse range of applications.
@@ III-B Selected LLMs
We evaluated a range of leading proprietary and open-source LLMs for their capacity to generate high-performance computing (HPC) code, including GPT-4o, Claude 3.5 Sonnet, DeepSeek Coder-V2, and Llama 3.1 8b. GPT-4o [openai_gpt4o] and Claude [anthropic_claude] are proprietary models renowned for their robust infrastructure, offering high-throughput performance and accessible APIs for integration into external systems. In contrast, DeepSeek Coder-V2 [deepseek_coder] and Llama 3.1 8b [llama3_2024] are prominent open-source models that can be deployed locally, enabling extensive customization and fine-grained control over parameters, memory allocation, and inference settings.
For the implementation of MARCO, we selected the ChatGPT (OpenAI) and Claude (Anthropic) APIs as primary LLM providers due to their mature developer ecosystems, comprehensive performance telemetry available through their developer consoles, and seamless integration with auxiliary technologies required in multi-agent workflows.
@@ III-C Benchmark Design
To rigorously assess MARCO’s optimization capabilities, we designed a benchmark suite comprising tasks that reflect core challenges in high-performance computing. These tasks span multiple categories:
  • Parallelization: We tested MARCO’s ability to generate and optimize code using OpenMP, CUDA, MPI, and Python’s asyncio, enabling concurrency both within single-threaded contexts and across multi-threaded or multi-GPU environments. This approach prevents excessive thread creation, which can degrade performance in scenarios with small workloads [ding2023hpc, nichols2024large].
  • Memory Optimization: We evaluated MARCO’s use of vectorization and cache utilization strategies to improve data locality and reduce memory access bottlenecks. Notably, the system leverages its integrated web-search component to retrieve cutting-edge optimization techniques from research publications indexed in IEEE, ACM, and arXiv, addressing the knowledge gap in pretrained LLMs [nichols2023hpc].
  • Algorithmic Efficiency: We examined classical algorithmic problems, including sorting and matrix multiplication, using MARCO’s testing agent to present detailed runtime and space complexity comparisons for each improvement over the baseline implementation. This enables users to make informed, data-driven decisions about which optimization strategies to adopt.

### v3

@@ I Introduction
High-performance computing (HPC) represents the pinnacle of computational power, utilizing clusters of computing resources to overcome the limitations of individual machines. HPC’s fundamental advantage lies in implementing parallel processing techniques that maximize processor cluster performance, enabling complex data processing and mathematical calculations that would otherwise be infeasible [Nichols_2024]. HPC has been instrumental in driving innovation across diverse domains including climate modeling, astrophysics simulations, pharmaceutical research, energy optimization, financial risk analysis, and training state-of-the-art machine learning models, particularly Large Language Models [chen2024landscapechallengeshpcresearch, ding2023hpc, hoffmann2022training, jin2023evaluation, narayanan2021efficient, yin2023ion].
@@
Large Language Models (LLMs) have revolutionized technology across multiple domains, democratizing access to sophisticated programming through natural language interfaces. These models have evolved from conversational assistants to autonomous systems capable of executing traditionally manual tasks, such as code generation [jiang2024survey, wang2024llm, liu2024evaluating, phuong2024TDD, li2024using]. The ability of LLMs to generate, test, and evaluate code has significantly enhanced software engineering capabilities, reducing development cycles and increasing productivity.
@@
However, despite their general utility, LLMs face significant challenges when applied to domain-specific knowledge requirements. According to the ParEval benchmark, both open-source and commercial LLMs demonstrate substantially lower accuracy in generating correct and efficient parallel code compared to serial code [Nichols_2024]. This limitation stems from LLMs’ training methodology, which typically involves diverse internet-sourced data rather than specialized domain knowledge. The relative scarcity of HPC code examples available online further exacerbates this knowledge gap [Nichols_2024, chen2024landscapechallengeshpcresearch]. While LLMs have also been successfully deployed in parallel program verification [aaron_iwomp24, sollenberger2024llm4vvexploringllmasajudgevalidation], the unique demands of high-performance computing—requiring rigorously tested and optimized code for parallelism, memory efficiency, and execution speed—raise a critical research question: How can we effectively leverage LLMs to address HPC-specific programming challenges comprehensively?
@@
A promising approach to addressing LLM limitations is the implementation of Artificial Intelligence (AI) Agents—autonomous LLM-based systems capable of performing complex tasks through customized workflows and specialized tools [xi2023survey, acharya2024agentic, li2024survey, gupta2024survey, sudalaimuthu2024enhancing]. AI agents extend beyond natural language processing to encompass decision-making, problem-solving, and execution capabilities. By utilizing external tools, these agents can access up-to-date information and effectively complete domain-specific tasks. Multi-agent systems further enhance these capabilities by distributing complex tasks across specialized agents, enabling the handling of work that exceeds the capabilities of any single agent.
@@ I-A Research Contributions
  • Specialized Agent Architecture: MARCO introduces a clear separation between code generation, performance evaluation, and testing agents, allowing each to develop distinct expertise. The code generation agent specializes in implementing advanced optimization techniques including cache locality optimization, parallelization through OpenMP or CUDA, and vectorization strategies. Meanwhile, the evaluation agent focuses on rigorously assessing execution time, memory usage, and algorithmic complexity, while the testing agent continuously evaluates output correctness.
@@
To assess the effectiveness of our approach, we first assembled a comprehensive evaluation dataset, which included the LeetCode 75 problem set along with an additional 10 challenging problems. Next, we implemented MARCO and compared its performance to leading language models such as GPT-4o, Claude 3.5 Sonnet, DeepSeek Coder-V2, and Llama 3.1. Additionally, we conducted detailed ablation studies to measure the contribution of each component within the MARCO architecture.
@@
The remainder of this paper is organized as follows: Section II discusses background and related work, Section III details our methodology, Section IV presents experimental results, Section V provides an in-depth analysis of the MARCO system, Section VI discusses our findings, and Section VII concludes with future research directions.
@@ II-A LLMs for Code Generation
LLMs can perform a wide range of tasks, such as code generation and understanding, through their parameters and transformer architecture [wang2024llm, liu2024evaluating, phuong2024TDD, li2024survey]. LLMs contain billions of parameters that enable them to identify patterns in the data encountered during training. Research demonstrates a strong correlation between parameter size and model effectiveness; as the number of parameters increases, the model’s ability to solve higher-complexity or data-intensive problems generally improves [kaplan2020scaling, hoffmann2022training, earle2023level]. The transformer architecture, a specialized type of neural network, is designed to process sequential data and predict intended outcomes through attention mechanisms.
@@
Chatbots like ChatGPT were among the first to provide general-purpose assistance across numerous domains. LLMs acquire domain knowledge through training on large datasets spanning diverse fields. OpenAI, for example, trained ChatGPT using techniques like Reinforcement Learning from Human Feedback [ouyang2022training, christiano2017deep, liu2024best], which utilizes a reward system to reinforce positive responses based on human trainer feedback. Given the adaptability of LLMs, specialization has emerged through the development of domain-specific models designed for particular applications. Code generation represents one popular specialization domain, where models are pre-trained primarily on code repositories to enhance their programming capabilities. Advanced code-focused LLMs like Qwen2.5-Coder employ techniques such as synthetic data generation to address the scarcity of high-quality code examples [hui2024qwen, wang2024llm].
@@
While LLMs have demonstrated impressive performance in code generation tasks, their effectiveness is primarily limited to small-scale problems, such as individual functions and single-file programs [li2024performance]. As coding problems increase in complexity and scale, these models experience significant performance degradation, largely due to limitations in their context processing capabilities [wang2024large]. This degradation relates directly to the LLM’s context window—the maximum amount of information the model can consider when generating responses, measured in tokens. LLMs process natural language by segmenting it into tokens, which can range from individual characters to complete words, enabling them to interpret and generate human-like text [balfroid2024towards]. Research has demonstrated that context window size critically determines an LLM’s effectiveness for complex tasks like code generation, as larger windows enable models to process more documentation and maintain awareness of broader code structures [antropic2023introducing]. This capability has evolved significantly, progressing from earlier models like GPT-3.5 with a context window of 16,385 tokens to more advanced models such as GPT-4 Turbo, which supports substantially expanded context windows [li2024performance].
@@ II-B Challenges in LLM-Generated HPC Code
LLM-generated code often exhibits significant deficiencies in performance optimization, including inefficient cache utilization and suboptimal parallelization strategies [nichols2024large]. The ParEval benchmark has revealed a substantial gap between LLMs’ ability to generate serial versus parallel code, with all evaluated models demonstrating markedly lower accuracy when generating parallel implementations [nichols2024large, davis2024pareval]. While open-source repositories provide vast amounts of training data for general coding tasks, this breadth comes at the expense of the specialized domain knowledge required for high-performance computing applications [ding2023hpc]. The primary challenge stems from how LLMs process information through Natural Language Processing (NLP) techniques, which excel with textual inputs but struggle with the computationally intensive scientific data prevalent in HPC contexts [chen2023lm4hpc].
Modern HPC infrastructure leverages parallel execution on multi-core systems through threading and synchronization mechanisms to optimize performance. However, due to the relatively recent emergence of these architectures, LLMs have limited understanding of modern parallel programming paradigms [kadosh2023scope]. This knowledge gap results in parallel code generation that performs substantially worse than sequential implementations [nichols2024large]. Furthermore, the advanced algorithms and optimization techniques typical in HPC environments create additional complexity that makes error diagnosis and correction particularly challenging for current LLM implementations [ding2023hpc].
These contextually demanding factors cause general-purpose LLMs to struggle with comprehensive HPC optimization. Recent research has begun addressing this problem with the introduction of HPC-Coder, which fine-tunes pre-existing models on HPC-specific data to improve parallel code generation capabilities [nichols2023hpc]. This specialized model outperforms general-purpose LLMs on HPC-specific tasks and OpenMP pragma generation, demonstrating the potential for targeted HPC optimization in LLM applications.
LLM-generated code also struggles with complex problems on initial attempts, often requiring multiple iterations to achieve correctness [huang2023agentcoder]. LLMs are inherently limited by single-pass generation approaches, which may be insufficient for challenging programming tasks. An agent-based approach offers a promising alternative for complex code generation, allowing multiple attempts at problem-solving and iterative code improvement [dong2023selfcollab]. This mirrors human problem-solving behavior, where programmers identify errors, learn from mistakes, and refine their solutions [yuqian2024colearning]. Given the limitations of LLMs in HPC code generation, simple single-model approaches may be inadequate for addressing these specialized requirements [singh2023selfplanning].
@@ III-A Problem Statement and Approach
Current approaches to code generation using Large Language Models (LLMs) exhibit significant limitations when applied to high-performance computing (HPC) workloads. Most existing models generate code in a single-pass manner, focusing primarily on syntactic correctness rather than optimizing for parallelism, memory hierarchy, or target machine characteristics. As a result, generated code often underperforms in memory efficiency, execution speed, and scalability, especially compared to hand-optimized or domain-specific implementations. This performance gap requires substantial manual tuning and expert intervention, highlighting the inadequacy of current LLM-driven workflows for HPC applications. Recent studies, including HPC-GPT [ding2023hpc] and the ParEval benchmark [nichols2024large], have empirically demonstrated that LLMs struggle to deliver performant parallel code, revealing a clear need for more sophisticated, architecture-aware generation techniques. Furthermore, work by Nichols et al. [nichols2023hpc] shows that even fine-tuned models on HPC-specific data require additional optimization layers to approach expert-written code performance.
@@
To address these challenges, we introduce MARCO (Multi-Agent Reactive Code Optimizer), a multi-agent system that fundamentally reimagines the code optimization process by combining automated code generation with systematic performance testing through API integrations for dynamic adaptation. This modular separation between generation and evaluation enables improvements beyond conventional monolithic approaches. MARCO’s architecture centers on four key innovations:
@@
  • Specialized Agent Separation: MARCO implements a clear division of responsibilities between the code optimizer and testing agent. The optimizer focuses on applying advanced strategies such as cache locality improvements, parallelization through OpenMP or CUDA, and vectorization techniques, drawing upon insights from recent high-performance computing literature [ding2023hpc, nichols2023hpc]. Meanwhile, the testing agent rigorously evaluates execution time, memory consumption, algorithmic complexity, output correctness, and incremental improvements over the original unoptimized code.
@@
  • Adaptive Feedback Loop: Unlike traditional single-pass LLM systems that seek immediate solutions, MARCO employs an iterative feedback mechanism. Performance metrics gathered by the testing agent are fed back to the optimizer, creating a refinement loop that progressively drives the system toward more efficient implementations, aligning with multi-step agentic workflows recently explored in AI systems research [huang2023agentcoder].
@@
  • Cost-Efficient Design: By decoupling optimization and evaluation tasks, MARCO reduces unnecessary API calls and minimizes token consumption at each iteration. This architectural design ensures controlled system costs, offering a scalable, resource-aware alternative to computationally expensive fine-tuning approaches.
@@
  • Broad Applicability: While many existing frameworks rely on specialized fine-tuning for narrow HPC tasks, MARCO dynamically adapts to live performance data and integrates real-time insights from web-sourced research. This generalizable design reduces the time and resources required for model retraining, maintaining efficiency across diverse applications.
@@ III-B Selected LLMs
We evaluated a range of leading proprietary and open-source LLMs for their capacity to generate high-performance computing code, including GPT-4o, Claude 3.5 Sonnet, DeepSeek Coder-V2, and Llama 3.1 8b. GPT-4o [openai_gpt4o] and Claude [anthropic_claude] are proprietary models renowned for their robust infrastructure, offering high-throughput performance and accessible APIs for integration into external systems. In contrast, DeepSeek Coder-V2 [deepseek_coder] and Llama 3.1 8b [llama3_2024] are prominent open-source models that enable local deployment, extensive customization, and fine-grained control over parameters, memory allocation, and inference settings.
For MARCO’s implementation, we selected the ChatGPT (OpenAI) and Claude (Anthropic) APIs as primary LLM providers due to their mature developer ecosystems, comprehensive performance telemetry available through developer consoles, and seamless integration with auxiliary technologies required in multi-agent workflows.
@@ III-C Benchmark Design
To rigorously assess MARCO’s optimization capabilities, we designed a benchmark suite comprising tasks that reflect core challenges in high-performance computing. These tasks span three primary categories:
  • Parallelization: We tested MARCO’s ability to generate and optimize code using OpenMP, CUDA, MPI, and Python’s asyncio, enabling concurrency within single-threaded contexts and across multi-threaded or multi-GPU environments. Our evaluation framework carefully manages thread creation to avoid performance degradation in scenarios with small workloads, a common pitfall in parallel code generation [ding2023hpc, nichols2024large].
  • Memory Optimization: We evaluated MARCO’s implementation of vectorization and cache utilization strategies to improve data locality and reduce memory access bottlenecks. The system leverages its integrated web-search component to retrieve cutting-edge optimization techniques from research publications indexed in IEEE, ACM, and arXiv, addressing the knowledge gap in pretrained LLMs [nichols2023hpc].
  • Algorithmic Efficiency: We examined classical algorithmic problems, including sorting and matrix multiplication, using MARCO’s testing agent to provide detailed runtime and space complexity comparisons for each improvement over baseline implementations. This enables users to make informed, data-driven decisions about optimization strategy adoption.